OpenStudy (anonymous):

The population of a town is 10,000 people. it grows at a rate of 1.4% per year for four years. what is the growth factor, b, for the equation y= a x b^x ?

OpenStudy (anonymous):

equation y= a x b^x = 10000 (1+1.4%)^4 the growth factor is (1+1.4%) = 1.014